# Learning Confident Classifiers in the Presence of Label Noise

### v2

@@ Abstract
The success of Deep Neural Network (DNN) models significantly depends on the quality of provided annotations. In medical image segmentation, for example, having multiple expert annotations for each data point is common to minimize subjective annotation bias. Then, the goal of estimation is to filter out the label noise and recover the ground-truth masks, which are not explicitly given. This paper proposes a probabilistic model for noisy observations that allows us to build a confident classification and segmentation models. To accomplish it, we explicitly model label noise and introduce a new information-based regularization that pushes the network to recover the ground-truth labels. In addition, for segmentation task we adjust the loss function by prioritizing learning in high-confidence regions where all the annotators agree on labeling. We evaluate the proposed method on a series of classification tasks such as noisy versions of MNIST, CIFAR-10, Fashion-MNIST datasets as well as CIFAR-10N, which is real-world dataset with noisy human annotations. Additionally, for segmentation task, we consider several medical imaging datasets, such as, LIDC and RIGA that reflect real-world inter-variability among multiple annotators. Our experiments show that our algorithm outperforms state-of-the-art solutions for the considered classification and segmentation problems.
@@ 1 Introduction
Real world data are replete with noisy labels. Since the labeling process of large-scale datasets is costly and time-consuming, researchers often resort to less expensive options, such as internet inquiries and crowdsourcing to circumvent this issue [53, 45]. Unfortunately, these methods are viable in producing datasets with incorrect labels. Smaller datasets are also vulnerable to the presence of corrupted labels. In this case, usually the labelling process is either challenging or the annotators have divergent opinions [4, 33]. In medical imaging, for example, it is imperative to procure annotations from the clinical experts. However, it is not only expensive to obtain annotated data, but it also suffers from high inter-reader variability among domain’s experts [26, 31], particularly in tasks such as segmentation and classification of images [26, 21]. Furthermore, the quality of the annotated labels became more intrinsic with the ubiquitous use of Deep Learning (DL) models in the medical domain [42, 49, 7], as the decision of the trainable models might have consequences for the health or even the life of a patient. Hence, developing DL algorithms that are robust to the noise in the annotated data is inarguably important.
DNNs noticeably suffer a degeneration in performance when trained on noisy labels. To combat this issue, various algorithms have been devised to adapt to the presence of noisy labels without compromising on the performance of DNNs. In classification, Sample Selection methods [14, 55, 34, 47, 19] started to gain momentum recently; these methods involve a two networks approach, Student-Teacher, for learning from noisy labels. It uses a small loss trick to sample clean instances for additional training by its peer network. However, it does not perform well when the loss distribution of true-labelled and false-labelled examples overlap substantially.
In segmentation, and particularly in the medical imaging domain, while automated and semi-automated algorithms are being developed, manual segmentation remains the gold standard for the quantitative interpretation of medical images and validation of those semi- or fully-automated algorithms. However, manual delineation by radiologists or other experts often relies on their subjective interpretation that is augmented by their expertise.
@@
The differences in annotations by multiple raters lead to inter-rater variability, which reflects the level of consistency between different raters in measuring the same target [21, 41]. This variability arises due to differences in experience, subject-specific expertise, and interpretation among the annotators. Moreover, the annotation task could be complex, time-consuming, and likely to be taxing; this inevitably leads to noise in the annotations, which degrades the model’s performance. To counter this, ground truth annotations are typically obtained through STAPLE [46] majority voting technique from multiple raters. This is done by selecting one preferred expert, or by using other label fusion approaches [23, 17, 3, 20, 57, 9, 27, 32, 37]. However, this process losses important agreement/disagreement information from all experts. In addition, while this technique is simple and easy to adopt, it comes at the expense of ignoring the underlying uncertainty information among various experts.
@@
In this work, we aim to develop a new deep learning approach that achieves high-quality predictions in the presence of multiple annotators in the dataset. Inspired by [44, 56], our approach considers coupling two convolutional neural network architectures: a base classification network that estimates the true labels and an annotation network that estimates confusion matrices for each rater for a given input image. However, unlike other methods, we want to make a base network confident in predicting each class in the presence of the noisy labels alone. To accomplish it, we introduce the new regularization term that forces predictions to be close to the one-hot vector, with $1$ being for the correct class and 00 for the rest. Such an approach assumes low ambiguity in ground truth annotations and attributes existing label uncertainty to the annotator’s noise. In addition, for segmentation tasks, we make use of the fact there are regions where all annotators agree and feed that information as input to the additional loss term that forces the confident predictions by the base segmentation network without the involvement of the annotation network.
Contributions. We summarize our contributions as follows.
@@
    We introduce a new “confidence” regularizer that allows to jointly learn the parameters of both base classification and annotation networks. It explicitly forces the model to learn confident predictions, thus allowing it to filter out the annotator noise in the problems with low ambiguity of ground truth labels.
@@
    For segmentation problem, we introduce a new term in the loss function, which takes into account the “agreement” between annotators in the regions where all of them give the same labels, thus forcing the direct application of the base segmentation model in the regions with a low annotation noise.
@@
    We provide a comprehensive set of experiments on MNIST, CIFAR-10, Fashion-MNIST and CIFAR-10N (real-world noisy data) datasets for classification. We also conducted a series of experiments on segmentation problems with annotation such as synthetic noisy data (based on MNIST) and real-world medical data (RIGA and LIDC). The experiments show that the proposed method provides state-of-the-art results in the considered problems.
@@ 2 Related Work
Learning with noisy labelled training data has been an active area of research for some time. Various algorithms have been introduced and have shown resistance to noise during training. We highlight the core research being done in this domain.
@@ 2.1 Classification
Noise Transition Matrix/Loss Correction. The loss correction approach [10, 6, 50] using a noise transition matrix, $T$, is a crucial branch that is used in deep learning systems. The goal of loss correction is for training on noisy labels with the corrected loss to be roughly equivalent to training on clean labels with the original loss.
@@
Patrini et al. [38] introduced two different approaches for loss correction using a stochastic matrix $T$ that delineates the probability of a class being flipped with another under a certain noise. To address extreme noise, Hendrycks et al. [15] suggested Gold Loss Correction (GLC) based on Forward Correction. Other similar works exploring this approach is [39, 12].
Multi-Network Learning. Multi-network [34, 19] training frequently employs collaborative learning and co-training. Therefore, the sample selection procedure is governed by the mentor network in the case of joint learning and the peer network in the case of co-training.
@@
Co-teaching [14] and Co-teaching+ [55] also employ two DNNs, but each DNN selects a certain number of small-loss examples and feeds them to its peer DNN for additional training. In comparison, JoCoR [47] reduces the diversity of two networks by means of co-regularization, making the predictions of the two networks more similar.
@@
Robust Regularization. Tanno et al. [44] showcased a method for simultaneously learning the individual annotator model and the underlying true label distribution through the use of trace regularization. Menon et al. [35] suggests a composite loss-based gradient clipping for label noise robustness, given that excessive trust is not placed in any single sample.
@@
Other Deep Learning/Statistical Methods. Yi and Wu [54] proposed a framework that updates the label distributions through an iterative correction of the noisy labels. Xia et al. [51] suggested a robust early-training method to diminish the side effect of noisy labels prior to early stopping. DivideMix [28] splits the data into labelled (clean samples) and an unlabelled set (noisy samples) and train it in a semi-supervised approach.
2.2 Segmentation
@@
Training deep neural networks on noisy label annotations can overfit to the noise in the dataset [43]. Inter-reader variability among annotators gave prominence to the Simultaneous Truth and Performance Level Estimation (STAPLE; [46]) algorithm that uses Expectation-Maximization to merge segmentation masks from various annotators into estimating a single ground truth. Several algorithms drew their inspiration from the STAPLE framework such as [23, 17, 3, 20, 57, 9, 27, 32, 37].
@@
In [36], authors provide a sample re-weighting strategy that considers the expertise level of annotators. This strategy gives greater weights in the loss function for the samples annotated by professionals. To disengage annotator bias, Zhang et al. [56] uses two coupled CNNs. Similar to [44], the CNN for segmentation estimates the label distribution, while the CNN for annotation is representative of the annotator bias using a confusion matrix. Similar ideas were applied in [29], but with another regularization that led to the recovery of the ground truth confusion matrix and ground truth posterior $p(y\mid x)$ under the assumption of diagonal dominance for the confusion matrix.
@@
Recently, various projects have begun to investigate the impact of multi-rater labels using label sampling [16, 22] or multi-head [13] techniques. Models trained with multi-rater labels are believed to be better calibrated than those trained with traditional ground-truth labels, such as majority vote, which are prone to overconfidence [16, 22]. In another pivotal work in the domain, Kohl et al. [24] proposed to learn a distribution of probable segmentations from the given input image; it integrates a integrates U-Net [40] and conditional Variational Autoencoders (VAE) to effectively develop an extensive number of conceivable hypothesis/segmentation results.
@@ 3 Methodology
3.1 Noisy Labels in Deep Learning
@@
It is worth noting that the ground truth annotations ${y_{i}}$ are not available to us but assumed to exist. Thus, our goal is to build a model which will approximate the distribution of targets $p({y_{i}}\mid x_{i})$ for a given input solely from the dataset of noisy annotations.
3.2 Probabilistic Model for Noisy Observations and Architecture
@@
$p(\{{\tilde{y}^{r}}\}_{r\in{S(x)}}\mid x)={\prod}_{r\in{S(x)}}p({\tilde{y}^{r}}\mid x).$ (1)
@@
Since the annotators’ noise is dependent on the sample $x$, this allows us to model noisy label distribution as $p(\tilde{y}^{r}=j\mid y=i,x)=:u^{r}_{ji}(x)$. These noisy class conditional distributions can be considered as elements of a confusion matrix (CM), effectively modeling the noise introduced by the annotator. We denote by $U^{r}(x)$ a $C\times C$ confusion matrix $[U^{r}]_{j,i}(x)=u^{r}_{ji}(x)$. Now, the marginalization can be rewritten as follows: {EQA} p(~y^r∣x) &= ∑_c = 1^C p(y= c ∣x)    p(~y^r∣y= c, x)
= ∑_c = 1^C u^r_~y^rc(x) ⋅p(y = c ∣x). In this equation, we see that the predictive distribution for a noisy annotation is a mixture of class-conditional distributions of noisy annotations weighed by the ground truth label distributions.
Inspired by [56, 44], we use a two-component model to approximate the distributions in equation (3.2). The first component is the base prediction network, which solves the problem of interest. In the case of classification, we call it classification model, in the case of segmentation – segmentation model. The other component is an annotation network. The base prediction network is a convolutional neural network parameterized by $\theta$ that aims to approximate the ground truth probability distribution $p({y}\mid x)$. Hence, it transforms an input image $x$ to a probability map $\hat{p}_{\theta}(x)\in\mathbb{R}^{C}$ (for more details in the case of segmentation see Section A).
The annotation network, parameterized by $\psi$, provides an estimate of the CMs of the corresponding annotators as a function of the input image. Hence, this model, trained separately for each of the annotators, produces the matrix $\hat{T}^{r}_{\psi}(x)\in[0,1]^{C\times C}$, $r\in{S(x)}$. The element of this matrix, indexed by $kl$, equals to $p({\tilde{y}^{r}}=k\mid{y}=l,x)$.
@@
4 Training and Confident Regularization
In this section, we explain in detail the training procedure we use and introduce a new regularizer specially tailored to the problems with label noise.
4.1 Loss function
@@
$-\log{\prod_{i=1}^{N}}{\prod_{r\in{S(x_{i})}}}\text{Cat}({\tilde{y}_{i}^{r}};p_{\psi,\theta}^{r}(x_{i}))=\\ -{\sum_{i=1}^{N}}{\sum_{r\in{S(x_{i})}}}\sum_{c=1}^{C}\mathbbm{1}({\tilde{y}_{i}^{r}}==c)\log p_{\psi,\theta}^{r}(x_{i})=\\ {\sum_{i=1}^{N}}{\sum_{r\in{S(x_{i})}}}CE\bigl{(}\hat{T}^{r}_{\psi}(x_{i})\,\hat{p}_{\theta}(x_{i}),{\tilde{y}_{i}^{r}}\bigr{)},$ (2)
where $CE$ stands for the cross-entropy and Cat stands for the probability mass function of Categorical distribution.
Minimizing the above objective function would jointly optimize both the parameters $\theta$ and $\psi$. However, minimizing this function alone does not guarantee disentangling the annotator’s noise distribution from the ground truth label distribution. This is because there could be many combinations of $\hat{T}^{r}_{\psi}(x)\,\hat{p}_{\theta}(x)$ that provide good approximations to ${\tilde{y}^{r}}$. This means that the prediction network’s output $\hat{p}_{\theta}(x)$ may not learn to correctly discriminate between noise and true labels. To overcome this issue, one may consider introducing some regularizer $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}$ that would help the disentanglement.
@@
$\mathcal{L}(X,Y)={\sum_{i=1}^{N}}{\sum_{r\in{S(x_{i})}}}CE\bigl{(}\hat{T}^{r}_{\psi}(x_{i})\,\hat{p}_{\theta}(x_{i}),{\tilde{y}_{i}^{r}}\bigr{)}+\\ +\lambda{\sum_{i=1}^{N}}R\bigl{(}\hat{p}_{\theta}(x_{i})\bigr{)}.$ (3)
In the next section we will discuss the particular choice of the regularizer $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}$.
4.2 Rationale for Confident Regularization
@@
where $\frac{1}{|{S(x)}|}\sum_{r\in{S(x)}}\hat{T}^{r}_{\psi}(x)$ is the confusion matrix averaged over annotators. The motivation of this regularizer is to push the estimated confusion matrix $\hat{T}^{r}_{\psi}(x)$ to converge to the true confusion matrix of the annotator. It is shown in [56], that the minimizer of the trace includes true confusion matrix as a solution.
@@
In this work, we propose a new regularizer that leads to better results without the need for regularizing parameters of the annotation network and the use of architecture modifications. Specifically, we push the base prediction network to provide confident predictions. Implicitly, we assume that there is no aleatoric noise in the ground-truth distribution of labels, the assumption previously considered in [56]. Hence, $p(y\mid x)=\textbf{e}$, where e is a one-hot vector with one placed to the correct label location. Based on these considerations, we introduce so-called confidence regularizer to the model’s loss function that penalizes the base prediction model for making uncertain predictions:
$R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}=\min\limits_{i}\bigl{(}-[\log\hat{p}_{\theta}(x)]_{i}\bigr{)}.$ (4)
This function attains minimum if one of the predicted classes has predicted probability equal to $1$.

### v3

@@ Abstract
The success of Deep Neural Network (DNN) models significantly depends on the quality of provided annotations. In medical image segmentation, for example, having multiple expert annotations for each data point is standard to minimize subjective annotation bias. Then, the goal of estimation is to filter out the label noise and recover the ground-truth masks, which are not explicitly given. This paper proposes a probabilistic model for noisy observations that allows us to build confident classification and segmentation models. We explicitly model label noise to accomplish this and introduce a new information-based regularization that pushes the network to recover the ground-truth labels. In addition, we adjust the loss function for the segmentation task by prioritizing learning in high-confidence regions where all the annotators agree on labeling. We evaluate the proposed method on a series of classification tasks such as noisy versions of MNIST, CIFAR-10, and Fashion-MNIST datasets, as well as CIFAR-10N, a real-world dataset with noisy human annotations. Additionally, for the segmentation task, we consider several medical imaging datasets, such as LIDC and RIGA, that reflect real-world inter-variability among multiple annotators. Our experiments show that our algorithm outperforms state-of-the-art solutions for the considered classification and segmentation problems.
@@ 1 Introduction
Real-world data are replete with noisy labels. Since the labeling process of large-scale datasets is costly and time-consuming, researchers often resort to less expensive options, such as internet inquiries and crowdsourcing to circumvent this issue [51, 44]. Unfortunately, these methods are viable in producing datasets with incorrect labels. Smaller datasets are also vulnerable to the presence of corrupted labels. In this case, usually, the labeling process is either challenging or the annotators have divergent opinions [4, 31]. In medical imaging, for example, it is imperative to procure annotations from clinical experts. However, it is not only expensive to obtain annotated data, but it also suffers from high inter-reader variability among domain’s experts [26, 30], particularly in tasks such as segmentation and classification of images [26, 20]. Furthermore, the quality of the annotated labels became more intrinsic with the ubiquitous use of Deep Learning (DL) models in the medical domain [40, 48, 7], as the decision of the trainable models might have consequences for the health or even the life of a patient. Hence, developing DL algorithms that are robust to the noise in the annotated data is inarguably important.
DNNs perform worse when there are noisy labels. Some algorithms have been created to help DNNs keep their performance even with noise. For example, Sample Selection methods that employ a Student-Teacher dual network approach have been developed to handle noisy labels, utilizing “small loss criteria” to identify clean instances for peer training, as discussed in works like [13, 53, 32, 46, 18]. However, their effectiveness decreases when true and false-labeled examples’ loss distributions significantly overlap.
@@
The differences in annotations by multiple raters lead to inter-rater variability, which reflects the level of consistency between different raters in measuring the same target [20, 39]. This variability arises due to differences in experience, subject-specific expertise, and interpretation among the annotators. Moreover, the annotation task could be complex, time-consuming, and likely to be taxing; this inevitably leads to noise in the annotations, which degrades the model’s performance. To counter this, ground-truth annotations are typically obtained through Simultaneous Truth and Performance Level Estimation (STAPLE; [45]) majority voting technique from multiple raters. This is done by selecting one preferred expert, or by using other label fusion approaches [23, 16, 3, 19, 55, 35]. However, this process loses important agreement/disagreement information from all experts. In addition, while this technique is simple and easy to adopt, it comes at the expense of ignoring the underlying uncertainty among various experts.
@@
In this work, we develop a new deep learning approach that achieves high-quality predictions in the presence of multiple annotators, that induce label noise in the dataset. This approach works both, for classification and for segmentation. Inspired by [43, 54], we consider coupling two models: a base classification network that estimates the true labels and an annotation network that estimates confusion matrices for each rater for a given input image. However, unlike other methods, we want to make a base network confident in predicting each class in the presence of the noisy labels alone. To accomplish it, we introduce the new regularization term that forces predictions to be close to the one-hot vector, with $1$ being for the correct class and 00 for the rest; Figure 1 gives an illustration of the details of the proposed architecture. This approach assumes low ambiguity in ground truth annotations and attributes existing label uncertainty to the annotator’s noise. In addition, for segmentation tasks, we make use of the fact there are regions where all annotators agree and feed that information as input to the additional loss term that forces the confident predictions by the base segmentation network without the involvement of the annotation network; details of the segmentation network is depicted in Figure 5.
We summarize our contributions as follows.
@@
    We introduce a class of “confidence” (information-based) regularizers, that allows us to jointly learn the parameters of both base classification and annotation networks. It explicitly forces the model to learn confident predictions, thus allowing it to filter out the annotator noise in the problems with low ambiguity of ground truth labels.
@@
    For the segmentation problem, we introduce a new term in the loss function, which takes into account the “agreement” between annotators in the regions where all of them give the same labels. This lets us directly use the base segmentation model in areas where there is little annotation noise.
@@
    We provide a comprehensive set of experiments on MNIST, CIFAR-10, Fashion-MNIST, and CIFAR-10N (real-world noisy data) datasets for classification. We also conducted a series of experiments on segmentation problems with annotation such as synthetic noisy data (based on MNIST) and real-world medical data (RIGA and LIDC). The experiments show that the proposed method provides state-of-the-art results in the considered problems.
@@ 2 Related Work
Learning with noisy labeled training data has been an active area of research for some time. Various algorithms have been introduced and have shown resistance to noise during training. We highlight the core research being done in this domain.
@@ 2.1 Classification
Noise Transition Matrix/Loss Correction. The loss correction approach [9, 6] using a noise transition matrix, $T$, is a crucial branch that is used in deep learning systems. The goal of loss correction is for training on noisy labels with the corrected loss to be roughly equivalent to training on clean labels with the original loss. The paper [36] introduced two different approaches for loss correction using a stochastic matrix $T$ that delineates the probability of a class being flipped with another under a certain noise. To address extreme noise, [14] suggested Gold Loss Correction (GLC) based on Forward Correction [42]. Other similar works exploring this approach are [37, 11].
@@
Multi-Network Learning. Multi-network [32, 18] training frequently employs collaborative learning and co-training. Therefore, the sample selection procedure is governed by the mentor network in the case of joint learning and the peer network in the case of co-training. Co-teaching [13] and Co-teaching+ [53] also employ two DNNs, but each DNN selects a certain number of small-loss examples and feeds them to its peer DNN for additional training. In comparison, JoCoR [46] reduces the diversity of two networks by means of co-regularization, making the predictions of the two networks more similar.
@@
Robust Regularization. In [43] showcased a method for simultaneously learning the individual annotator model and the underlying true label distribution through the use of trace regularization. The paper [33] suggests a composite loss-based gradient clipping for label noise robustness, given that excessive trust is not placed in any single sample.
@@
Other Deep Learning/Statistical Methods. The authors of [52] proposed a framework that updates label distributions by repeatedly correcting the noisy labels. The paper [49] suggested a robust early-training method to diminish the side effects of noisy labels prior to early stopping. DivideMix [27] splits the data into labeled (clean samples) and an unlabelled set (noisy samples) and train it in a semi-supervised approach.
2.2 Segmentation.
@@
Training deep neural networks on noisy label annotations can overfit the noise in the dataset [41]. Inter-reader variability among annotators gave prominence to the STAPLE; [45]) algorithm that uses Expectation-Maximization to merge segmentation masks from various annotators into estimating a single ground truth. Several algorithms drew their inspiration from the STAPLE framework such as [23, 16, 3].
@@
In [34], authors provide a sample re-weighting strategy that considers the expertise level of annotators. This strategy gives greater weights in the loss function for the samples annotated by professionals. To disengage annotator bias, the paper [54] uses two coupled DNNs. Similar to [43], the DNN for segmentation estimates the label distribution, while the DNN for annotation is representative of the annotator bias using a confusion matrix. Similar ideas were applied in [28], but with another regularization that led to the recovery of the ground truth confusion matrix and ground truth posterior $p(y\mid x)$ under the assumption of diagonal dominance for the confusion matrix.
@@
Recently, various projects have begun to investigate the impact of multi-rater labels using label sampling [15, 21] or multi-head [12] techniques. Models trained with multi-rater labels are believed to be better calibrated than those trained with traditional ground-truth labels, such as majority vote, which are prone to overconfidence [15, 21]. The paper [24] proposed learning a distribution of possible segmentations from a given input image. It combines U-Net [38] with conditional Variational Autoencoders (VAE) to generate a wide range of potential segmentation outcomes.
@@ 3 Methodology
3.1 Noisy Labels in Deep Learning.
@@
It is worth noting that the ground truth annotations ${y_{i}}$ are not available to us but assumed to exist. Thus, our goal is to build a model that will approximate the distribution of targets $p({y_{i}}\mid x_{i})$ for a given input solely from the dataset of noisy annotations.
3.2 Probabilistic Model for Noisy Observations and Architecture.
@@
(3.1) $p(\{{\tilde{y}^{r}}\}_{r\in{S(x)}}\mid x)={\prod}_{r\in{S(x)}}\,p({\tilde{y}^{r}}\mid x).$
@@
Since the annotators’ noise is dependent on the sample $x$, this allows us to model noisy label distribution as $p(\tilde{y}^{r}=j\mid y=k,x)=:u^{r}_{jk}(x)$. These noisy class conditional distributions can be considered as elements of a confusion matrix (CM), effectively modeling the noise introduced by the annotator.
Now, the marginalization can be rewritten as follows:
(3.2) $p({\tilde{y}^{r}}\mid x)=\sum\nolimits_{c=1}^{C}u^{r}_{{\tilde{y}^{r}}c}(x)\cdot p(y=c\mid x).$
In this equation, we see that the predictive distribution for a noisy annotation is a mixture of class-conditional distributions of noisy annotations weighed by the ground truth label distributions.
Inspired by [54, 43], we use a two-component model to approximate the distributions in equation (3.2). The first component is the base prediction network, which solves the problem of interest. In the case of classification, we call it the classification model, in the case of segmentation – segmentation model. The other component is an annotation network. The base prediction network is a convolutional neural network parameterized by $\theta$ that aims to approximate the ground truth probability distribution $p({y}\mid x)$. Hence, it transforms an input image $x$ to a probability map $\hat{p}_{\theta}(x)\in\mathbb{R}^{C}$ (for more details in the case of segmentation see Section A).
The annotation network, parameterized by $\psi$, is designed to estimate the confusion matrices associated with each annotator, treating them as functions of the input image. Specifically, for each annotator $r$, the network produces a matrix $\hat{T}^{\psi}_{r}(x)\in[0,1]^{C\times C},$ where $C$ is the number of classes. Each entry in this matrix, denoted as $\hat{T}^{\psi}_{r}(x)[k,l]$, represents the estimated probability $p({\tilde{y}^{r}}=k\mid{y}=l,x)$ that an annotator labels a sample of true class $l$ as class $k$, conditioned on the input $x$.
@@
4 Training and Confidence Regularization
In this section, we explain in detail the training procedure we use and introduce a new class of regularizers, specially tailored to the problems with label noise.
4.1 Loss function.
@@ 4 Training and Confidence Regularization
(4.3) $-\log{\prod_{i=1}^{N}}{\prod_{r\in{S(x_{i})}}}\text{Cat}({\tilde{y}_{i}^{r}};p_{\psi,\theta}^{r}(x_{i}))\\ =\textstyle\sum_{i=1}^{N}\sum_{r\in{S(x_{i})}}CE\bigl{(}\hat{T}^{r}_{\psi}(x_{i})\,\hat{p}_{\theta}(x_{i}),{\tilde{y}_{i}^{r}}\bigr{)},$
where $CE$ means the cross-entropy and Cat is the probability mass function of Categorical distribution.
Minimizing the above objective function would jointly optimize both the parameters $\theta$ and $\psi$. However, minimizing this function alone does not guarantee to disentangle the annotator’s noise distribution from the ground truth label distribution. This is because there could be many combinations of $\hat{T}^{r}_{\psi}(x)\,\hat{p}_{\theta}(x)$ that provide good approximations to ${\tilde{y}^{r}}$. This means that the prediction network’s output $\hat{p}_{\theta}(x)$ may not learn to correctly discriminate between noise and true labels. To overcome this, one may consider introducing some regularizer $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}$ that would help the disentanglement.
@@
$\sum_{i=1}^{N}{\sum_{r\in{S(x_{i})}}}CE\bigl{(}\hat{T}^{r}_{\psi}(x_{i})\,\hat{p}_{\theta}(x_{i}),{\tilde{y}_{i}^{r}}\bigr{)}+\lambda\sum_{i=1}^{N}R\bigl{(}\hat{p}_{\theta}(x_{i})\bigr{)}.$
In the next section, we will discuss the particular choice of the regularizer $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}$.
4.2 Rationale for Confidence Regularization.
@@
where $\frac{1}{|{S(x)}|}\sum_{r\in{S(x)}}\hat{T}^{r}_{\psi}(x)$ is the confusion matrix averaged over annotators. The motivation of this regularizer is to push the estimated confusion matrix $\hat{T}^{r}_{\psi}(x)$ to converge to the true confusion matrix of the annotator. It is shown in [54], that the minimizer of the trace includes a true confusion matrix as a solution.
@@
In this work, we propose a new class of regularizers, that leads to better results without the need for regularizing parameters of the annotation network and the use of architecture modifications. Specifically, we push the base prediction network to provide confident predictions. Implicitly, we assume that there is no aleatoric noise in the ground-truth distribution of labels, the assumption previously considered in [54]. This assumption, despite seeming restrictive, holds in many image analysis datasets [22] and can be reached in practice, given better instuctions to annotators. Hence, $p(y\mid x)=\textbf{e}$, where e is a one-hot vector with one placed to the correct label location. Based on these considerations, we propose a class of regularizes, that encourages base model to provide confident predictions. We consider two instances of this class: the so-called confidence regularizer:
(4.4) $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}=\min\limits_{c\in\overline{1,C}}\bigl{(}-[\log\hat{p}_{\theta}(x)]_{c}\bigr{)}.$
This function attains minimum if one of the predicted classes has a predicted probability equal to $1$.
As another instance of this class, we consider predictive entropy regularizer:
(4.5) $R\bigl{(}\hat{p}_{\theta}(x)\bigr{)}=-\sum\nolimits_{c=1}^{C}\bigl{[}\hat{p}_{\theta}(x)\bigr{]}_{c}\bigl{[}\log\hat{p}_{\theta}(x)]_{c}\bigr{.}$